OpenStudy (anonymous):

Simplify this expression. 4^6 · 8^2

OpenStudy (anonymous):

First convert this problem into a^2 - b^2 format

OpenStudy (anonymous):

Whoa. Wait how do i do that?

OpenStudy (anonymous):

4^6 = 4*4*4*4*4*4 = 64*64 = 64^2 or, 4^6 = {4^3}^2 = 64^2

OpenStudy (anonymous):

then ur equation will get reduced to 4^6 - 8^2 = 64^2 - 8^2 that is in this format a^2 - b^2 the formula for this one is a^2-b^2 = (a-b)*(a+b)

OpenStudy (anonymous):

so now u can do this one

OpenStudy (anonymous):

I'll show you how i figured the answer out and you tell me if it's right and if there's any easier way

OpenStudy (anonymous):

4^6 x 8^2= (2^2)^6 x (2^3)^2= 2^(2x6) x 2^(3x2)= 2^12 x 2^6 = 2^(12 + 6) = 2^18 This could be a bit hard to follow

OpenStudy (anonymous):

oh god there was dot my mistake 4^6 X 8^2 = 2^12 X 2^6 = 2^18

OpenStudy (anonymous):

How do you simplify it??? It's confusing me cuz i simply guessed then did the math and even then i was confused

OpenStudy (anonymous):

4^6 = {2^2}^6 = 2^12 8^2 = {2^3}^2 = 2^6 2^12 X 2^6 = 2^{12+6} = 2^18

OpenStudy (anonymous):

4 = 2 X 2 = 2^2 8 = 2 X 2 X 2 = 2^3

OpenStudy (anonymous):

any other problem disturbing my dear friend

OpenStudy (anonymous):

Yeah... Psychic huh? :D

OpenStudy (anonymous):

hahahaha...................that was hilarious

OpenStudy (anonymous):

the medal was for Psychic one

OpenStudy (anonymous):

hahaha... thanks! aright next question (in case you haven't already guessed, i suck at math) :P

OpenStudy (anonymous):

(5^4)^6/5^8 Step by step?

OpenStudy (anonymous):

ok first take the numerator (5^4)^6 = 5^(4 X 6) = ???

OpenStudy (anonymous):

5^24?

OpenStudy (anonymous):

right now 5^24 / 5^8 = 5^(24-8) = ???

OpenStudy (anonymous):

5^16

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok then buddy bed calling nice to meet you.......will catch u later

OpenStudy (anonymous):

OK nice meeting you!